OpenStudy (anonymous):

Using the formula A=2h(L+W) find L if A=729.8, H=17.8, and W=6.4

OpenStudy (anonymous):

Hi Hannah! This time its a case of both substitution and rearranging. So first lets rearrange like before to fine L.

OpenStudy (anonymous):

A=2h(L+W) find L if A=729.8, H=17.8, and W=6.4 A = 2hL + 2hW A - 2hW = 2hL (A - 2hW)/(2h) = L (729.8 - ((2)(17.8)(6.4)))/(2*17.8) = L 501.96/35.6 = L L = 14.1

OpenStudy (anonymous):

You do realise you have to let some people do their own learning? You just have to prod them down the right path?